You are considering investing ​$72,000 in new equipment. You estimate that the net cash flows will be ​$11,00 during the first​ year, but will increase by ​$1,700 per year the next year and each year thereafter. The equipment is estimated to have a 7​-year service life and a net salvage value of ​$6,000 at that time. Assume an interest rate of 9​%.

​(a) Determine the annual capital cost​ (ownership cost) for the equipment. The annual capital cost is ​$______. ​(Round to the nearest​ dollar.)

(b) Determine the equivalent annual savings (revenues). The equivalent annual savings are $________. (Round to the nearest dollar.)

(c) Determine whether this is a wise investment. The annual equivalent worth is $______. (Round to the nearest dollar.)

Is this a good investment?

yes or no
Answer #1

​(a) Annual capital cost​ for the equipment = 72000(A/P, 9%, 7) = $14,306

(b) Equivalent annual savings = 1,100 + 1,700(A/G, 9%, 7) + 6000(A/F, 9%, 7) = $6,270

(c) The annual equivalent worth is $6,270 - $14,306 = -$8,036

Since the annual worth is a negative value, this is not a good investment.